# Scaling Large Motion Models with Million-Level Human Motions

### v2

@@ Abstract
Inspired by the recent success of LLMs, the field of human motion understanding has increasingly shifted toward developing large motion models. Despite some progress, current efforts remain far from achieving truly generalist models, primarily due to the lack of massive high-quality data. To address this gap, we present MotionLib, the first million-level dataset for motion generation, which is at least 15$\times$ larger than existing counterparts and enriched with hierarchical text descriptions. Using MotionLib, we train a large motion model named Being-M0, demonstrating robust performance across a wide range of human activities, including unseen ones. Through systematic investigation, for the first time, we highlight the importance of scaling both data and model size for advancing motion generation, along with key insights to achieve this goal. To better integrate the motion modality, we propose Motionbook, an innovative motion encoding approach including (1) a compact yet lossless feature to represent motions; (2) a novel 2D lookup-free motion tokenizer that preserves fine-grained motion details while expanding codebook capacity, significantly enhancing the representational power of motion tokens. We believe this work lays the groundwork for developing more versatile and powerful motion generation models in the future. For further details, visit our project page.
@@ 1 Introduction
Figure 1: TOP: While existing models perform well on Motion-X and HumanML3D, they struggle with out-of-domain concepts on MotionLib, exhibiting limited generalization. DOWN: Curves showing the effects of scaling up large motion models. MotionLib is the first large T2M dataset comparable in scale to visual benchmarks like ImageNet.
@@ 2 Related Work
Large Language Models and Multi-Modality. Substantial advancements have been made in enhancing LLMs (Brown et al., 2020; Raffel et al., 2020; Chowdhery et al., 2022) with the ability to understand and respond to human instructions, through a technique known as instruction tuning (Ouyang et al., 2022). Recent research has extended these capabilities to the multimodal domain (Ye et al., 2023), with notable work by Liu et al. (2023), who pioneered visual instruction tuning to create a highly adaptable visual assistant. Additionally, Li et al. (2023a) integrated multimodal context directly into instruction data to further enhance performance. Subsequent studies (Zhang et al., 2023b) expand this research by scaling up instructional datasets and incorporating image-rich text. Notably, Dai et al. (2023) developed InstructBLIP, based on BLIP-2 (Li et al., 2023b), which features an advanced visual feature extraction mechanism to improve performance across vision-language tasks. Despite these breakthroughs, the application of multimodal models to human motion remains less competitive compared to current SoTA methods, although recent initiatives are beginning to explore this domain (Zhang et al., 2024b).
@@
Human Motion Generation The task of motion generation involves creating human motion based on various inputs, such as text descriptions (Guo et al., 2022b), action labels (Cervantes et al., 2022) or motion prefixes (Liu et al., 2022). Among these, text-to-motion (T2M) generation has received the most attention due to the ease and flexibility of using natural language as input. Early approaches (Fragkiadaki et al., 2015; Ghosh et al., 2017) rely on deterministic motion modeling, which often produces averaged, blurry results. To overcome this, stochastic methods using models like GANs (Wang et al., 2020) or VAEs (Aliakbarian et al., 2020) have been considered. For instance, T2M-GPT (Zhang et al., 2023a) extends the temporal VAE to capture the probabilistic relationship between text and motion. Recently, Guo et al. (2024) proposed integrating residual quantization and masked modeling to improve traditional vector quantization (VQ). Lu et al. (2023) designed HumanTomato, a hierarchical VQ-VAE to separately encode body and hand motions. To better align with a motion auto-encoder, MotionCLIP (Tevet et al., 2022) incorporates CLIP (Radford et al., 2021) as the text encoder, bringing in more robust text priors. Additionally, Zhang et al. (2024b) and Jiang et al. (2023) explored the development of unified models based on LLMs that accept multimodal conditions (e.g., vision, text, and pose), enabling the generation of subsequent, preceding, or “in-between” motions. Despite leveraging the power of LLMs, these large motion models remain limited to in-domain text instructions and do not yet perform as competitively as existing SoTA methods.
@@ 3 MotionLib: A Million-Level Motion Library
Data are the foundation of large motion models. With advancements in fields such as human pose detection, we are now able to extract high-quality motion sequences from vast amounts of online videos. Our MotionLib contains over one million motion clips, each annotated with fine-grained automatic pseudo-labels. A comparison with existing benchmarks is presented in Table 1. Examples of this dataset are shown in Figure 2. Our data curation pipeline involves the following main procedural steps. For more details, please refer to Appendix B due to limited space.
@@
Hierarchical Motion Descriptions. Existing benchmarks face inherent limitations in their text descriptions. Previous studies (Guo et al., 2022a) typically use a few short sentences to describe whole-body motions, neglecting finer details of individual body parts, such as the arms or legs. This restricts the model from performing more nuanced body comprehension and more flexible part-level motion control (e.g., raising only the left arm). Moreover, the richness of text labels often varies across different motions; for example, a large portion of the Motion-X dataset provides only a single sentence. In contrast, MotionLib offers hierarchical textual annotations for each video inspired by  Pi et al. (2023): (1) part-level description, where we carefully design a prompt format and use Gemini-1.5-pro (Reid et al., 2024) to generate detailed captions for individual body parts (e.g., left arm), assigning a dedicated sentence to each; (2) body-level description, which summarizes the whole body movement in a comprehensive paragraph containing 1–3 sentences; The hierarchical motion captions provide 2$\times$ more textual contents, which we believe is essential for the LLM to better align with the motion modality.
Motion and Description Refinement. It is inevitable that the million-scale motions collected above are noisy. According to some previous studies (Holden, 2024), the low-quality motion may damage generation results. To enhance the quality of our collected motion, we first estimate precise 3D keypoints with another pre-trained model (Sárándi et al., 2023) to perform local-global motion optimization, following Lin et al. (2024). In addition, we train an RL-based policy $\pi_{\text{refine}}$ based on  Luo et al. (2023) to elaborate and refine raw motions to obey physical laws (e.g., maintaining balance) and appear more realistic. $\pi_{\text{refine}}$ takes raw motion sequences as input, treating them as target poses, and generates new motion sequences that satisfy physical laws in a simulated environment, thus eliminating issues like jittering and foot-sliding. While effective, $\pi_{\text{refine}}$ may struggle with drastic movements in target poses, leading to slipping within the simulation. For such cases, we mark them with a smaller weight during pretraining, forcing our model to focus more on high-quality motion. To refine the description, since our text labels are structured into body and part levels, we condition GPT-4o with one level of text while using the other as the target to be refined. GPT-4o then assesses and refines the text content of the target level, enabling the generation of more precise and reliable descriptions.
@@ 4.1 Overview
Similar to previous LLM-based multimodal models, we treat motion as a foreign language. The overall framework is illustrated in Figure 3. Our large motion model, built on a pre-trained LLM, functions as a generative model that connects a motion tokenizer with the LLM backbone $\Theta$. The motion tokenizer encodes the features =$\mathcal{M}=\{m_{1},m_{2},...,m_{T}\}$ of raw motion sequence into token embeddings $\mathcal{V}=\{v_{1},v_{2},...,v_{n}\}\in\mathbbm{R}^{n\times d}$, where $n$ denotes the number of motion tokens and $d$ represents the dimensionality of each token. To integrate motion tokens into the LLM backbone, we incorporate $K$ discrete codes in the motion codebook as additional vocabulary for the LLM. Additionally, we introduce two special tokens, $<$mot$>$ and $<$/mot$>$, to signify the start and end of motion sequences within the input/output streams. The LLM backbone $\Theta$ is built on a decoder-only architecture using causal transformers. The model generates outputs $\mathcal{Y}=\{y_{1},y_{2},...,y_{m}\}$ in an auto-regressive manner, where $\mathcal{Y}$ corresponds to the generated motion sequence based on the provided motion-text input tokens. In this work, each motion-text pair in the MotionLib dataset is framed as an instruction-following instance $\{\mathcal{X}_{Q},\mathcal{X}_{M}\}$, representing a question-answer interaction between the user and the motion model. The entire instructional dataset adheres to this unified format. To train our model, we optimize the negative log-likelihood over the predicted tokens which is defined as:
@@
Two-Stage Training. Similar to Liu et al. (2023), our training process incorporates motion-text alignment and motion instruction tuning. In the first stage, we utilize the entire MotionLib dataset. In the second stage, to better align with real-world human commands, we construct a set of over 250 instruction templates (e.g., ”Show me a demonstration of $<$Caption_Placeholder$>$ through movement.”). We then select a subset of high-quality motions from MotionLib to formulate our instructional data. For each motion, we randomly choose three templates, inserting the corresponding text into the $<$Caption_Placeholder$>$. To further enhance naturalness and fluency, we leverage Gemini-Pro to refine these instructions, ensuring they align with human expression patterns. Finally, we construct an instruction fine-tuning dataset containing 900K instructions.
@@ 4.2 MotionBook: Towards Effective Motion Encoding
MotionBook consists of a lossless feature to represent motion sequence $\mathcal{M}$ and a 2D-LFQ motion tokenizer to compress $\mathcal{M}$ into a series of discrete token.
@@
In addition, previous motion tokenizers are restricted from capturing the full diversity of human motion due to their small codebook sizes. A common approach to address this limitation is to expand the vocabulary. However, excessively enlarging the codebook can result in codebook collapse, where only a small subset of tokens in the codebook is used, offering minimal improvements even potentially degrading the model’s performance. A more effective way is to reduce the dimensionality of code embeddings (Mentzer et al., 2023), which limits the representational capacity of individual tokens and encourages more efficient learning across a larger vocabulary. Similar to Yu et al. (2023), we reduce the embedding dimension of the codebook to zero by replacing the codebook $\mathbbm{C}\in\mathcal{R}^{K\times d}$ with an integer set $\mathbbm{C}$ with $|\mathbbm{C}|=K$. Here, $\mathbbm{C}$ is the Cartesian product of single-dimensional variables $\mathbbm{C}=$$\times$ ${}_{i=1}^{d}C_{i}$, where $C_{i}=\{-1,1\}$ and $d$ is equal to $\log_{2}K$. Given a feature vector $z\in\mathbbm{R}^{d}$, our motion quantizer $Q(\cdot)$ converts each dimension of the quantized representation into:
@@ 5.1 Experimental Setup
Evaluation and Implementation Details. For the motion generation task, we employ the evaluation metrics following (Guo et al., 2022a): Motion-retrieval Precision (R-Precision), Multimodal Distance (MMDist), and Frechet Inception Distance (FID). In addition, we evaluate our motion tokenizer on the motion reconstruction task.
@@ 5.2 Discussion of Scaling up Motion Generation
In this section, we investigate the impact of model size and data scale for motion generation. We utilize the motion autoencoder (Guo et al., 2022a) retrained on Motion-X and MotionLib to evaluate performance on their respective test sets. We categorize training data into four scales: 0.02M (HumanML3D only), 0.08M (Motion-X only), 0.5M (MotionLib-0.5), and 1.2M (MotionLib-full). For fair comparison, we employ the same motion tokenizer, maintaining consistency across experiments to validate our conclusions.
Does increasing model size benefit motion generation? Yes. As shown in Table 2, the experimental results demonstrate that increasing model size leads to a stable performance enhancement when the training data is kept constant. Specifically, the performance of LLaMA2-13b surpasses that of LLaMA2-7b, which in turn outperforms GPT-2. This series of results strongly validates the trend of performance gains scaling with model capacity. Consequently, larger models show a greater ability to capture the diverse and intricate patterns and relationships within human motion.
Does increasing the data scale benefit motion generation? Yes. As shown Table 2, when using the same foundation model, increasing the scale of training data leads to substantial improvement on MotionLib testing set, aligning with our expected scaling laws. This improvement is particularly pronounced in the R-precision metric, emphasizing the critical role of data scale in enhancing semantic alignment between generated motions and text prompts. It’s worth noting that after pretraining on MotionLib, the performance of models trained on Motion-X also improves. This demonstrates that large-scale pretraining allows the model to have a better initialization for motion-language alignment.
@@
Does our large motion model outperform SoTA models? Yes. We evaluate our model Being-M0 on the widely adopted HumanML3D benchmark. We compare its performance against a variety of SoTA approaches. This includes diffusion-based methods such as MLD (Chen et al., 2023) and MotionDiffuse (Zhang et al., 2022), as well as the GPT-based T2M-GPT (Zhang et al., 2023a). We also compare against LLM fine-tuning methods like MotionGPT (Jiang et al., 2023; Zhang et al., 2024b), MotionGPT-v2 (Wang et al., 2024), MotionLLM (Wu et al., 2024), and AvatarGPT (Zhou et al., 2024). As shown in Table 3, our model, which utilizes LLaMA2-13B as the motion decoder, achieves SOTA performance. Our model significantly outperforms other LLM-based methods such as AvatarGPT and the latest MotionGPT-v2. In particular, we observe substantial improvements in key metrics such as R@1, R@3, and MMDist, highlighting our model’s ability to generate motion sequences that are better aligned with text descriptions and of higher quality.
@@
Does large motion model excel in out-of-distribution setup? Yes. We present the results in Table LABEL:tab:out_of_distribute. This ablation is essential for further validating the true generalization capabilities of large motion models, as the improvements observed in Table 2 may result from the inclusion of additional in-domain data similar to testbeds. In this setup, we select four subsets from MotionLib, comprising 90K samples (UNSEEN-90K), for evaluation, while the remaining 38 subsets are used for training. This ensures that the testing set consists entirely of out-of-domain (OOD) samples. We compare the performance of models trained on HumanML3D, Motion-X, and Motion-#11, all utilizing the GPT-2 architecture, where #$\#N$ denotes the number of training subsets. The results on the OOD test set clearly demonstrate that the model trained on MotionLib significantly outperforms those trained on HumanML3D and Motion-X, particularly in terms of R@1 and R@3 metrics. These findings strongly highlight the superior generalization ability of large motion models when handling unseen OOD data, especially when trained on diverse, large-scale datasets.
@@
Does motion instruction tuning help? We compare experiments with and without instruction tuning on Motion-X. As shown in Table LABEL:tab:inst_tune, instruction tuning data effectively improves various metrics. This indicates a benefit for the model’s further language-motion alignment capability. Instruction tuning also makes the model more user-friendly, proving it a worthwhile optimization.

### v3

@@ Abstract
Inspired by the recent success of LLMs, the field of human motion understanding has increasingly shifted toward developing large motion models. Despite some progress, current efforts remain far from achieving truly generalist models, primarily due to the lack of massive high-quality data. To address this gap, we present MotionLib, the first million-level dataset for motion generation, which is at least 15$\times$ larger than existing counterparts and enriched with hierarchical text descriptions. Using MotionLib, we train a large motion model named Being-M0, demonstrating robust performance across a wide range of human activities, including unseen ones. Through systematic investigation, for the first time, we highlight the importance of scaling both data and model size for advancing motion generation, along with key insights to achieve this goal. To better integrate the motion modality, we propose Motionbook, an innovative motion encoding approach including (1) a compact yet lossless feature to represent motions; (2) a novel 2D lookup-free motion tokenizer that preserves fine-grained motion details while expanding codebook capacity, significantly enhancing the representational power of motion tokens. We believe this work lays the groundwork for developing more versatile and powerful motion generation models in the future. For further details, visit https://beingbeyond.github.io/Being-M0/.
@@
Figure 1: TOP: While existing models perform well on small-scale datasets like Motion-X and HumanML3D, they struggle with out-of-domain concepts on MotionLib, exhibiting limited generalization. DOWN: Curves showing the effects of scaling up large motion models. MotionLib is the first large T2M dataset comparable in scale to visual benchmarks like ImageNet.
@@ 2 Related Work
Large Language Models and Multi-Modality. Substantial progress has been made in enhancing LLMs (Brown et al., 2020; Raffel et al., 2020; Chowdhery et al., 2022) with the ability to understand and respond to human commands, through a technique known as instruction tuning (Ouyang et al., 2022). Recent research has extended these capabilities to the multimodal domain (Ye et al., 2023; Zheng et al., 2024; Feng et al., 2024; Mei et al., 2024), with notable work by Liu et al. (2023), who pioneered visual instruction tuning to create a highly adaptable assistant. In addition, Li et al. (2023a) integrated multimodal context into instruction data to further enhance performance. Subsequent studies (Zhang et al., 2023c) expanded this research by scaling up instructional data and incorporating image-rich text. Notably, Dai et al. (2023) developed InstructBLIP based on BLIP-2 (Li et al., 2023b), which features an advanced visual feature extraction mechanism to improve performance across vision-language tasks. Despite these, the application of multimodal models to human motion remains less competitive compared to current SoTA methods, although recent initiatives are beginning to explore this domain (Zhang et al., 2024d).
@@
Human Motion Generation The task of motion generation involves creating human motion based on various inputs, such as text descriptions (Guo et al., 2022b), action labels (Cervantes et al., 2022) or motion prefixes (Liu et al., 2022). Among these, text-to-motion (T2M) generation has received the most attention due to the ease and flexibility of using natural language as input. Early approaches (Fragkiadaki et al., 2015; Ghosh et al., 2017) rely on deterministic motion modeling, which often produces averaged, blurry results. To overcome this, stochastic methods using models like GANs (Wang et al., 2020) or VAEs (Aliakbarian et al., 2020) have been considered. For instance, T2M-GPT (Zhang et al., 2023a) extends the temporal VAE to capture the probabilistic relationship between text and motion. Recently, Guo et al. (2024) proposed integrating residual quantization and masked modeling to improve traditional vector quantization (VQ). Lu et al. (2023) designed HumanTomato, a hierarchical VQ-VAE to separately encode body and hand motions. To better align with a motion auto-encoder, MotionCLIP (Tevet et al., 2022) incorporates CLIP (Radford et al., 2021) as the text encoder, bringing in more robust text priors. Additionally, Zhang et al. (2024d) and Jiang et al. (2023) explored the development of unified models based on LLMs that accept multimodal conditions (e.g., vision, text, and pose), enabling the generation of subsequent, preceding, or “in-between” motions. Despite leveraging the power of LLMs, these large motion models remain limited to in-domain text instructions and do not yet perform as competitively as existing SoTA methods.
@@ 3 MotionLib: A Million-Level Motion Library
Data are the foundation of large motion models. With advancements in fields such as human pose detection, we are now able to extract high-quality motion sequences from vast amounts of online videos. Our MotionLib dataset contains over one million motion clips, totaling approximately 137 million frames. Each clip is annotated with fine-grained automatic pseudo-labels. A comparison with existing benchmarks is presented in Table 1. Examples of this dataset are shown in Figure 2. Our data curation pipeline involves the following main procedural steps. For more details, please refer to Appendix B due to limited space.
@@
Hierarchical Motion Descriptions. Existing benchmarks face inherent limitations in their text descriptions. Previous studies (Guo et al., 2022a) typically use a few short sentences to describe whole-body motions, neglecting finer details of individual body parts, such as the arms or legs. This restricts the model from performing more nuanced body comprehension and more flexible part-level motion control (e.g., raising only the left arm). Moreover, the richness of text labels often varies across different motions. For example, a large portion of the Motion-X dataset provides only a single sentence, and HumanML3D contains numerous similar text annotations. In contrast, MotionLib offers hierarchical textual annotations for each video inspired by  Pi et al. (2023): (1) part-level description, where we carefully design a prompt format and use Gemini-1.5-pro (Reid et al., 2024) to generate detailed captions for individual body parts (e.g., left arm), assigning a dedicated sentence to each; (2) body-level description, which summarizes the whole body movement in a comprehensive paragraph containing 1–3 sentences. These hierarchical motion captions provide far more textual content and levels of detail than previously available. Their richness and structure are crucial for large language models to better understand and align with the motion modality.
Motion and Description Refinement. It is inevitable that the million-scale motions collected from diverse web sources initially contain noise. According to some previous studies (Holden, 2024), the low-quality motion may damage generation results. To enhance the quality of our collected motion, we first estimate precise 3D keypoints with another pre-trained model (Sárándi et al., 2023) to perform local-global motion optimization, following Lin et al. (2024). In addition, we train an RL-based policy $\pi_{\text{refine}}$ based on  Luo et al. (2023) to elaborate and refine raw motions to obey physical laws (e.g., maintaining balance) and appear more realistic. $\pi_{\text{refine}}$ takes raw motion sequences as input, treating them as target poses, and generates new motion sequences that satisfy physical laws in a simulated environment, thus eliminating issues like jittering and foot-sliding. While effective, $\pi_{\text{refine}}$ may struggle with drastic movements in target poses, leading to slipping within the simulation. For such cases, we mark them with a smaller weight during pretraining, forcing our model to focus more on high-quality motion. To refine the description, since our text labels are structured into body and part levels, we condition GPT-4o with one level of text while using the other as the target to be refined. GPT-4o then assesses and refines the text content of the target level, enabling the generation of more precise and reliable descriptions.
@@ 4.1 Overview
Inspired by the success of Large Language Models (LLMs) in multimodal tasks, we adapt a similar approach for motion generation, treating motion as a distinct “language” to be learned (Luo et al., 2020; Zhang et al., 2024c). Similar to previous LLM-based multimodal models, we treat motion as a foreign language. The overall framework is illustrated in Figure 3. Our large motion model, built on a pre-trained LLM, functions as a generative model that connects a motion tokenizer with the LLM backbone $\Theta$. The motion tokenizer encodes the features =$\mathcal{M}=\{m_{1},m_{2},...,m_{T}\}$ of raw motion sequence into token embeddings $\mathcal{V}=\{v_{1},v_{2},...,v_{n}\}\in\mathbbm{R}^{n\times d}$, where $n$ denotes the number of motion tokens and $d$ represents the dimensionality of each token. To integrate motion tokens into the LLM backbone, we incorporate $K$ discrete codes in the motion codebook as additional vocabulary for the LLM. Additionally, we introduce two special tokens, $<$mot$>$ and $<$/mot$>$, to signify the start and end of motion sequences within the input/output streams. The LLM backbone $\Theta$ is built on a decoder-only architecture using causal transformers. The model generates outputs $\mathcal{Y}=\{y_{1},y_{2},...,y_{m}\}$ in an auto-regressive manner, where $\mathcal{Y}$ corresponds to the generated motion sequence based on the provided motion-text input tokens. In this work, each motion-text pair in the MotionLib dataset is framed as an instruction-following instance $\{\mathcal{X}_{Q},\mathcal{X}_{M}\}$, representing a question-answer interaction between the user and the motion model. The entire instructional dataset adheres to this unified format. To train our model, we optimize the negative log-likelihood over the predicted tokens which is defined as:
@@
Two-Stage Training. Similar to Liu et al. (2023), our training process incorporates motion-text alignment and motion instruction tuning. In the first stage, we utilize the entire MotionLib dataset to enable the model to learn fundamental motion-text correlations from diverse data. In the second stage, to better align with real-world human commands, we construct a set of over 250 instruction templates (e.g., ”Show me a demonstration of $<$Caption_Placeholder$>$ through movement.”). We then select a subset of high-quality motions from MotionLib to formulate our instructional data. For each motion, we randomly choose three templates, inserting the corresponding text into the $<$Caption_Placeholder$>$. To further enhance naturalness and fluency, we leverage Gemini-Pro to refine these instructions, ensuring they align with human expression patterns. Finally, we construct an instruction fine-tuning dataset containing 900K instructions.
@@ 4.2 MotionBook: Towards Effective Motion Encoding
MotionBook consists of (1) a lossless feature representation for motion sequences $\mathcal{M}$, and (2) a 2D Lookup-Free Quantization (2D-LFQ) motion tokenizer to compress $\mathcal{M}$ into a series of discrete tokens.
@@
In addition, previous motion tokenizers are restricted from capturing the full diversity of human motion due to their small codebook sizes. A common approach to address this limitation is to expand the vocabulary. However, excessively enlarging the codebook can result in codebook collapse, where only a small subset of tokens in the codebook is used, offering minimal improvements even potentially degrading the model’s performance. To enable effective learning with larger codebooks suited for million-scale datasets like MotionLib, we adopt a lookup-free quantization strategy. A more effective way is to reduce the dimensionality of code embeddings (Mentzer et al., 2023), which limits the representational capacity of individual tokens and encourages more efficient learning across a larger vocabulary. Similar to Yu et al. (2023), we reduce the embedding dimension of the codebook to zero by replacing the codebook $\mathbbm{C}\in\mathcal{R}^{K\times d}$ with an integer set $\mathbbm{C}$ with $|\mathbbm{C}|=K$. Here, $\mathbbm{C}$ is the Cartesian product of single-dimensional variables $\mathbbm{C}=$$\times$ ${}_{i=1}^{d}C_{i}$, where $C_{i}=\{-1,1\}$ and $d$ is equal to $\log_{2}K$. Given a feature vector $z\in\mathbbm{R}^{d}$, our motion quantizer $Q(\cdot)$ converts each dimension of the quantized representation into:
@@ 5.1 Experimental Setup
Evaluation. For the motion generation task, we employ the evaluation metrics following (Guo et al., 2022a): Motion-retrieval Precision (R-Precision), Multimodal Distance (MMDist), and Frechet Inception Distance (FID). R-Precision assesses the consistency or matching between generated motions and input text descriptions. MMDist measures the average distance between generated motion and corresponding text in feature space. FID evaluates the realism and quality of generated motions by comparing the distribution of generated motion features with that of real motion features. In addition, we evaluate our motion tokenizer on the motion reconstruction task. Besides FID, this task is also measured by MPJPE, which measures the average distance between predicted and ground truth joint positions across all joints, in millimeters.
@@ 5.2 Discussion of Scaling up Motion Generation
In this section, we investigate the impact of model size and data scale for motion generation. For evaluating the models, we use the same motion autoencoder architecture (Guo et al., 2022a) retrained separately on Motion-X and MotionLib. This model is used to evaluate the performance of motion generation models on their respective test sets. We categorize training data into four scales: 0.02M (HumanML3D only), 0.08M (Motion-X only), 0.5M (MotionLib-0.5), and 1.2M (MotionLib-full). For fair comparison, we employ the same motion tokenizer, maintaining consistency across experiments to validate our conclusions.
Does increasing model size benefit motion generation? Yes. As shown in Table 2, the experimental results demonstrate that increasing model size leads to a stable performance enhancement when the training data is kept constant. For example, our model achieves best R@1 (0.491) on Motion-X when using LLaMA2-13b, which is 0.016 higher than LLaMA2-7b’s 0.475. On the larger MotionLib dataset, the best R@1 of LLaMA2-13b reaches 0.185, which is 0.014 higher than LLaMA2-7b’s 0.171. This series of results strongly validates the trend of performance gains scaling with model capacity. Consequently, larger models show a greater ability to capture the diverse and intricate patterns and relationships within human motion.
Does increasing the data scale benefit motion generation? Yes. As illustrated in Table 2, when using the same foundation model, increasing the scale of training data leads to substantial improvement on MotionLib testing set, aligning with our expected scaling laws. When using LLaMA2-13b as the LLM backbone for our model, training with 1.2M data achieves an R@1 of 0.185, which is 0.011 and 0.061 higher than performance using 0.5M data and 0.08M Motion-X, respectively. This improvement is particularly pronounced in the R-precision metric, emphasizing the critical role of data scale in enhancing semantic alignment between generated motions and text prompts. It’s worth to note that the performance of models trained on Motion-X also improves, after pretraining on MotionLib. This demonstrates that large-scale pretraining allows the model to have a better initialization for motion-language alignment.
@@
Does our large motion model outperform SoTA models? Yes. We evaluate our model Being-M0 on the widely adopted HumanML3D benchmark. We compare its performance against a variety of SoTA approaches. This includes diffusion-based methods such as MLD (Chen et al., 2023), MotionDiffuse (Zhang et al., 2022), ReMoDiffuse (Zhang et al., 2023b), Fg-T2M++ (Wang et al., 2025) and LMM (Zhang et al., 2024a). It also includes autoregressive models like T2M-GPT (Zhang et al., 2023a), DiverseMotion (Lou et al., 2023), and MoMask (Guo et al., 2024). We also compare against LLM fine-tuning methods like MotionGPT (Jiang et al., 2023; Zhang et al., 2024d), MotionGPT-v2 (Wang et al., 2024), MotionLLM (Wu et al., 2024), and AvatarGPT (Zhou et al., 2024). As shown in Table 3, our model, which utilizes LLaMA2-13B as the motion decoder, achieves SOTA R-Precision performance. Existing T2M methods can be categorized intow two types: (1) specialist models optimized on specific datasets, and (2) LLM-based generalist models (aiming for broader instruction and task generalization via LLMs), like AvatarGPT, MotionGPT-v2 and our Puppet. Puppet excels among generalist models and remains highly competitive on R@1, R@3, and MMDist compared to specialist models. This highlights our model’s ability to generate motion sequences that are better aligned with text descriptions and of higher quality.
@@
Does large motion model excel in out-of-distribution setup? Yes. We present the results in Table LABEL:tab:out_of_distribute. This ablation is essential for further validating the true generalization capabilities of large motion models, as the improvements observed in Table 2 may result from the inclusion of additional in-domain data similar to testbeds. In this setup, we select 11 subsets from MotionLib, totaling 90K samples (UNSEEN-90K), for evaluation. The remaining subsets are used for training. These 11 subsets include both synthetic and real-world data. They cover domains like fitness, person-object interaction, social activities, and diverse environments such as labs and outdoors. The training set, on the other hand, primarily consists of Motion-X and web-derived data. This ensures that the test set is entirely composed of out-of-domain (OOD) samples. We compare the performance of models trained on HumanML3D, Motion-X, and Motion-#11, all utilizing the GPT-2 architecture, where #$\#N$ denotes the number of training subsets. The results on the OOD test set clearly demonstrate that the model trained on MotionLib significantly outperforms those trained on HumanML3D and Motion-X, particularly in terms of R@1 and R@3 metrics. These findings strongly highlight the superior generalization ability of large motion models when handling unseen OOD data, especially when trained on diverse, large-scale datasets. This demonstrates the value of utilizing a large amount of web-based motion data.
@@
TRAIN TEXT R@1 $\uparrow$ R@3 $\uparrow$ FID $\downarrow$
Real 0.297 0.634 0.004
single-level 0.162 0.371 7.018
hierarchical 0.166 0.375 6.936
@@
Does motion instruction tuning help? Yes. We compare experiments with and without instruction tuning on Motion-X. As shown in Table LABEL:tab:inst_tune, instruction tuning data effectively improves various metrics, with R@1 increased by 0.017 and FID increased by 0.010. This indicates a benefit for the model’s further language-motion alignment capability. Instruction tuning also makes the model more user-friendly, proving it a worthwhile optimization.
Does richer text description help? Yes. To investigate the effectiveness of richer text descriptions, we compared the training results using hierarchical descriptions, which include both body-level and part-level labels, against those using only whole-body description. As shown in Table LABEL:tab:hiera_basic, training with hierarchical descriptions improves both R@1 and R@3 by 0.004 compared to single-level descriptions. These results indicate that hierarchical text significantly enhances the model’s semantic understanding, leading to improved semantic matching of generated motions.